OpenStudy (afloridagirl):

When 27 times x squared times z all over negative 3 times x squared times z to the sixth power is completely simplified, the exponent on the variable z is _____.

OpenStudy (afloridagirl):

i need help really bad, i think it is 3 but im not sure

OpenStudy (agent0smith):

Post a screenshot or use the equation editor.

OpenStudy (afloridagirl):

27x^2z/-3x^2z^6

OpenStudy (afloridagirl):

thats the equation @agent0smith

OpenStudy (agent0smith):

You gotta learn to use the equation editor... is it this \[\large \frac{ 27x^2z}{ -3x^2z^6 }\]

OpenStudy (afloridagirl):

yes

OpenStudy (agent0smith):

So just simplify everything using exponent rules

OpenStudy (afloridagirl):

i did and got three is that right?

OpenStudy (agent0smith):

Show work cos i have no idea what you mean by just "three"

OpenStudy (afloridagirl):

z's exponent is 3

OpenStudy (agent0smith):

Just show your whole work on the simplification, and it def isn't 3

OpenStudy (afloridagirl):

oh

OpenStudy (agent0smith):

\[\Large \frac{ x^a }{x^b } = x^{a-b}\]

OpenStudy (afloridagirl):

27×x2×z−3×x2×z4⇒−9×x2×z−3×x2×z3⇒−9z3⇒−9z−3

OpenStudy (afloridagirl):

thats what i got....

OpenStudy (agent0smith):

I can't follow that... time to start practicing the equation editor.

OpenStudy (afloridagirl):

im probably completely wrong

OpenStudy (afloridagirl):

can you help me solve it

OpenStudy (afloridagirl):

please

OpenStudy (agent0smith):

large \frac{ 27x^2z}{ -3x^2z^6 } copy and paste that into the equation editor and start simplifying it. I don't understand how you got an exponent of 3 on the z.

OpenStudy (afloridagirl):

im trying but it wont do anyhing

OpenStudy (afloridagirl):

is there a possibility that it could be five?

OpenStudy (agent0smith):

\[\large \frac{ 27x^2z}{ -3x^2z^6 }\] \[\large \frac{ -9x^2z}{ x^2z^6 }\] \[\large \frac{ -9z}{ z^6 }\] This is the correct step by step working so far. I can not follow yours, in your working, the exponent on the z seems to change randomly for no reason?

OpenStudy (afloridagirl):

im probably doing it wrong

OpenStudy (agent0smith):

Well you need to start using the equation editor so people can follow your work, the way you wrote it is not easy to follow

OpenStudy (afloridagirl):

im sorry

OpenStudy (agent0smith):

Now simplify the last step, using the exponent rule i posted earlier.

OpenStudy (afloridagirl):

would it be -3

OpenStudy (agent0smith):

Tell me exponent on the z on the top of the fraction.

OpenStudy (agent0smith):

Or tell me why you think it's -3?

OpenStudy (afloridagirl):

because you add the 6 to the -9

OpenStudy (afloridagirl):

im completely wrong arent i?

OpenStudy (agent0smith):

The -9 has absolutely nothing to do with the z's btw. It being there makes no difference.

OpenStudy (agent0smith):

\[\large \frac{ -9z^1}{ z^6 }\]use the exponent rule, remembering that the z on top has an exponent of 1. The -9 has no effect at all, ignore it.

OpenStudy (afloridagirl):

oh okay ...

OpenStudy (afloridagirl):

so we subtract 1 and six

OpenStudy (afloridagirl):

soo it would be five?

OpenStudy (afloridagirl):

i mean 6 and one

OpenStudy (afloridagirl):

sorry

OpenStudy (agent0smith):

do you think \[\large \frac{ -9z^1}{ z^6 } \] becomes \[\large -9z^5\] or \[\large \frac{ -9}{ z^5 }\] which do you think is correct?

OpenStudy (afloridagirl):

the first one? @agent0smith

OpenStudy (afloridagirl):

im sorry i was on the phone with a teacher, @agent0smith

undeadknight26 (undeadknight26):

Good Ol'dba. Did good did ya?

OpenStudy (leenathan):

you know if u had an example for this formula it would be much much easier.

OpenStudy (afloridagirl):

im sorry can someone just tell me the answer?

OpenStudy (legomyego180):

5

OpenStudy (afloridagirl):

are you sure?

OpenStudy (legomyego180):

If you rewrite it and take the z out of the denominator it would be -5

OpenStudy (afloridagirl):

so -5

OpenStudy (agent0smith):

\[\large \frac{ -9z^1}{ z^6 }\]simplifies to \[\large -9z^{-5}\] since 1-6 is -5. But you shouldn't finish a problem with negative exponents, and hopefully you remember that you can move a negative exponent into the denominator to make it positive, so \[\large \frac{ -9}{ z^5 }\]

OpenStudy (afloridagirl):

so 5

OpenStudy (legomyego180):

^ what he said. The answer is not -5 though.\[\frac{ 9z }{ z^6 } = \frac{ 9 }{ z^5 } = 9z ^{-5}\]

OpenStudy (afloridagirl):

it is 5 thank you @agent0smith